OpenStudy (anonymous):

Both triangles are similar. The area of the smaller triangle is about 97 ft^sqrt. Which is the best approximation for the area of the larger triangle? smaller triangle side: 12ft bigger triangle side: 21ft

OpenStudy (anonymous):

Could any of you give me an equation so i can solve the problem on my own?

OpenStudy (dangerousjesse):

Do you know the type of triangle..?

OpenStudy (anonymous):

Equilateral sides.

OpenStudy (dangerousjesse):

Oh! Okay

OpenStudy (tkhunny):

The type of triangle is not significant. Linear Measure Proportion a:b Area Proportion: a^2:b^2 We know the similar triangle sides have the proportion 12:21 ==> 4:7 Thus, the area of the similar triangles must have the proportion 4^2:7^2 ==> 16:49 That's about it. How do we get 97 ft^2 in there?

OpenStudy (anonymous):

Here's a picture of the shapes.

OpenStudy (dangerousjesse):

Knowing the type of triangle lets me use an easy equation.. Sorry?

OpenStudy (anonymous):

It's okay, just do whatever it takes you need to do! :)

OpenStudy (triciaal):

double check the side of the smaller and use that for the final calculation 12 vs 15

OpenStudy (triciaal):

area of similar triangles is in ratio of sides as explained by @tkhunny above the area of the smaller triangle is 97 so the area of the larger is 97 * (factor) where factor is the inverse ratio of the lengths used. if 15:21 then 97 * 7/5 if 12:21 then 3*4/3*7 then it would be 97 *7/4

OpenStudy (tkhunny):

Not Quite... Length Ratio 4:7 Area Ratio 16:49 \(\dfrac{Smaller\;Area}{Larger\;Area} = \dfrac{16}{49}\) \(\dfrac{97\;ft^{2}}{Larger\;Area} = \dfrac{16}{49}\) \(Larger\;Area = 97\;ft^2\cdot\dfrac{49}{16}\)